OpenStudy (whatever):

can somone helpme.ive been trying to do this but i havent understand how to 10x+5y=2250 x/10+y/5=400

OpenStudy (anonymous):

first off multiply the second by 10 to get rid of the stupid decimal

OpenStudy (anonymous):

now you can see plainly that the ratio of y to x is the same, so these are either the same line or two parallel lines

OpenStudy (anonymous):

dang i meant \[10x+5y=2250\] \[2x+y=800\]

OpenStudy (anonymous):

there that is better

OpenStudy (anonymous):

so because the ratio is the same, two x for every one y in both of these, they have the same slope. so there is either no solution or they are the same line

OpenStudy (anonymous):

if you multiply the second equation by 5 you get \[10x+5y=2250\] \[100x+5y=4000\] and now you can see that they are not the same line, they are parallel lines. so no solution to this system

OpenStudy (whatever):

oh okay thank you